OpenStudy (anonymous):

Suppose the amount of money in a bank account with an initial balance of $4,000 earns 1.3 percent interest per year. If the exponential function, , is used to model this situation, which value would represent B in the function? 4,000 1.13 1.013 1.3

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Looks the exponential function has gone while copy pasting

OpenStudy (anonymous):

oh its y=A * B^x

ganeshie8 (ganeshie8):

Thanks! what is 1.3% in decimal form ?

OpenStudy (anonymous):

1.3

ganeshie8 (ganeshie8):

Nope, you need to divide by 100 remember

OpenStudy (anonymous):

no wait. 0.013

ganeshie8 (ganeshie8):

Yes!

ganeshie8 (ganeshie8):

B = 1 + 0.013 = ?

OpenStudy (anonymous):

1.013

OpenStudy (anonymous):

so option c is my answer?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

Thanks!